# MagR: Weight Magnitude Reduction for Enhancing Post-Training Quantization

### v1

@@ 4.2 MagR via $\ell_{\infty}$-Regularization
where $\alpha>0$ serves as the regularization parameter, balancing fidelity against the $\ell_{\infty}$ regularizer. To maintain the output of the layer, $\alpha$ should typically be set to a small value.
@@
Figure 2: The $\ell_{1}$-ball projection in $\mathbb{R}^{2}$.
Fortunately, computing projection onto the $\ell_{1}$ ball is an established task, and there are several efficient algorithms available. For example, see condat2016fast and the references therein. Here we adopted a simple algorithm of $O(n\log n)$ time complexity as in duchi2008efficient , which supports parallelizable or vectorized implementation for the projections of a batch of weight vectors, i.e., a weight matrix, as will be described in the next subsection. The implementation mainly involves sorting and soft-thresholding yin2008bregman ; see Algorithm 3 in Appendix A.1 for the details.
@@
where the proximal operator $\mathrm{prox}_{t\|\cdot\|_{\infty}}$ and the corresponding projection $\mathrm{proj}_{\|\cdot\|_{1}\leq 1}$ in (4) are applied column-wise to the matrix input. Hereby we summarize MagR for processing one linear layer in Algorithm 1 with the column-wise $\ell_{1}$-ball projection for matrix input as detailed in Algorithm 2.
@@
Input: Pre-trained weight matrix $\hat{\boldsymbol{W}}\in\mathbb{R}^{m\times n}$; Hessian matrix $\boldsymbol{H}=\boldsymbol{X}^{\top}\boldsymbol{X}$; max iteration number $K$; step size $\eta=\frac{1}{\lambda_{\max}(\boldsymbol{H})}$; penalty parameter $\alpha>0$.
@@
MagR for Per-Group Quantization. By using more float scaling factors, per-group quantization becomes a preferred strategy for mitigating accuracy loss at extremely low bit-widths. In this approach, a weight vector $\mathbf{w}\in\mathbb{R}^{m}$ is segmented into groups of weights, each containing $d$ elements, with all weights within a group sharing a common scaling factor for quantization. Here, per-group MagR applies an $\ell_{\infty}$ penalty to each vector of grouped weights. Consequently, the $\ell_{1}$-ball projection is independently performed on these vectors, while maintaining the gradient descent step unchanged. We note that the group-wise $\ell_{1}$-ball projection can be easily done using Algorithm 2, with an additional reshaping of the input $\boldsymbol{V}\in\mathbb{R}^{m\times n}$ into $\mathbb{R}^{d\times\left(\frac{m}{d}\cdot n\right)}$.
@@ 5 Experiments
We employed our MagR processing approach on top of the two gradient-free PTQ methods, RTN and OPTQ frantar2022optq , to quantize the LLaMA1 (7B-65B) touvron2023llama and LLaMA2 (7B-70B) touvron2023llama2 model families. By applying MagR on top of RTN (MagR+RTN), we achieved better results than AWQ lin2023awq for per-channel INT3 and INT4 weight quantization. Additionally, MagR combined with OPTQ (MagR+OPTQ) achieved state-of-the-art performance for INT3 and INT4 qunatization. To enhance the per-channel INT2 quantization, we ran 30 additional iterations of coordinate descent algorithm zhang2024comq on top of OPTQ, which we denote by MagR+OPTQ. It turns out MagR+OPTQ is superior to both Omniquant shao2023omniquant and QuIP chee2024quip in terms of perplexity (Table 2), and falls just short of QuIP in zero-shot tasks for 13B and 70B models (Table 3). Note that QuIP uses random orthogonal transformations (so-called Incoherence Processing) to process both the weights and features, resulting in 1.5×1.5$\times$1.5 × slower throughput than OPTQ. In contrast, MagR-based method does not introduce any overhead whatsoever compared with OPTQ.
@@
The choice of parameters. To ensure that the MagR-processed layer output $\boldsymbol{X}\boldsymbol{W}$ is faithful to the original $\boldsymbol{X}\hat{\boldsymbol{W}}$, we need to use a tiny penalty parameter $\alpha$ in (2). For per-channel quantization, $\alpha$ was fixed to be $10^{-3}$ in our experiments, but we did find that setting it to a smaller value of $5\times 10^{-4}$ or $10^{-4}$ can sometimes slightly improve the perplexity (with a relative change of $<1\%$ in ppl). Similarly for per-group quantization, we set $\alpha$ to $10^{-4}$, while reducing it to $5\times 10^{-5}$ or $10^{-5}$ could sometimes also slightly improve the perplexity.
Furthermore, we used a multiplicative scalar $\beta<1$ to decay the standard quantization step $\delta=\frac{\max(\boldsymbol{w})-\min(\boldsymbol{w})}{2^{b}-1}$ (or equivalently, the quantization scale) of the quantizer. In other words, our $\delta=\beta\cdot\frac{\max(\boldsymbol{w})-\min(\boldsymbol{w})}{2^{b}-1}$. It has been shown in existing works li2016ternary ; rastegari2016xnor that, optimal quantization step for binary or ternary quantization yielding the minimum quantization error is not given by $\frac{\max(\boldsymbol{w})-\min(\boldsymbol{w})}{2^{b}-1}$. Shrinking $\delta$ at low bit-width results in a more clustered quantization grid lattice that fits the weights better, which leads to a smaller overall error. In general, $\beta$ is positively correlated with the bit-width used. For per-channel quantization, the best $\beta\in[0.8,0.85]$ on INT2 quantization, whereas the empirically optimal $\beta$ is around $0.9$ for INT3 quantization. As for INT4, $\beta$ is simply set to 1, that is, we used the standard quantization step. In addition, for per-group quantization, we chose $\beta=0.95$ for both INT2 and INT3 quantization. We observed that this refinement on the quantization step $\delta$ significantly improves the performance of the PTQ method. In addition, the iteration number $K$ in Algorithm 1 was set to $150$ across all the experiments.
@@ 5.1 Language Generation
We concentrate our analysis on perplexity-based tasks. The results for the LLaMA2 family with context length of 2048, are elaborated in Table 2, while those for LLaMA1 are provided in Appendix Table 5. As evidenced by the tables, the MagR preprocessing consistently improve the performance of the baselines RTN and OPTQ. Moreover, MagR+OPTQ consistently outperforms most baseline across the LLaMA family models for both per-channel and per-group weight quantization. Particularly, for INT2, MagR+OPTQ performs 30 additional coordinate descent (CD) iterations on top of OPTQ to refine the solution, surpassing all baselines.
@@ 5.2 Zero-Shot Tasks
We evaluated the performance of quantized models on several zero-shot tasks. The results are reported in Table 3. Similar to previous observations, the proposed MagR demonstrates superior performance on most models compared to OmniQuant, with a small gap compared to QuIP chee2024quip . Nonetheless, it is reasonable and commendable that our algorithm achieves results close to QuIP without introducing any inference overhead. It is possible to further improve our approach based on the insight behind QuIP chee2024quip - i.e., quantization benefits from incoherent weight and Hessian matrices. As shown in Table 3, incoherence processing often improves the quantization of large models, especially at higher compression rates (e.g., 2bits). The extension of MagR into incoherence preprocessing is left for future work.
Table 4: The runtime of MagR+RTN, MagR+OPTQ, and MagR+OPTQ, with comparisons to their vanilla counterparts, namely, RTN and OPTQ.
@@ Appendix A Appendix / supplemental material
A.1 Projection of a Vector onto $\ell_{1}$-Ball
Algorithm 3 describes the implementation of unit $\ell_{1}$-ball projection of a vector in $\mathbb{R}^{m}$.
Algorithm 3 Projection onto the unit $\ell_{1}$-ball.
@@ A.1 Projection of a Vector onto $\ell_{1}$-Ball
1:  Sort $|\boldsymbol{v}|$ into $\boldsymbol{\mu}$ such that $\mu_{1}\geq\mu_{2}\geq\ldots\geq\mu_{m}$.
2:  if $\|\boldsymbol{v}\|_{1}>\epsilon$ then

### v2

@@ 4.2 MagR via $\ell_{\infty}$-Regularization
where $\alpha>0$ serves as the regularization parameter, balancing fidelity against the $\ell_{\infty}$ regularizer. To maintain the output of the layer, $\alpha$ should typically be set to a small value. Indeed, let $\boldsymbol{w}^{*}$ be the minimizer of (2), we have that the $\ell_{2}$ error of the layer’s output introduced by MagR is $O(\sqrt{\alpha})$:
$\|\boldsymbol{X}\boldsymbol{w}^{*}-\boldsymbol{X}\hat{\boldsymbol{w}}\|$ $\leq\sqrt{\|\boldsymbol{X}\boldsymbol{w}^{*}-\boldsymbol{X}\hat{\boldsymbol{w}}\|^{2}+2\alpha\|\boldsymbol{w}^{*}\|_{\infty}}$
$\leq\sqrt{\|\boldsymbol{X}\hat{\boldsymbol{w}}-\boldsymbol{X}\hat{\boldsymbol{w}}\|^{2}+2\alpha\|\hat{\boldsymbol{w}}\|_{\infty}}=\sqrt{2\alpha\|\hat{\boldsymbol{w}}\|_{\infty}},$
where $\|\hat{\boldsymbol{w}}\|_{\infty}$ is a constant independent of $\alpha$, and the second inequality uses that $\boldsymbol{w}^{*}$ is the minimizer.
@@
Fortunately, computing projection onto the $\ell_{1}$ ball is an established task, and there are several efficient algorithms available. For example, see condat2016fast and the references therein. Here we adopted a simple algorithm of $O(m\log m)$ time complexity as in duchi2008efficient , which supports parallelizable or vectorized implementation for the projections of a batch of weight vectors, i.e., a weight matrix, as will be described in the next subsection. The implementation mainly involves sorting and soft-thresholding yin2008bregman ; see Algorithm 3 and its derivation in Appendix A.1 for the details.
@@
where the proximal operator $\mathrm{prox}_{t\|\cdot\|_{\infty}}$ and the corresponding projection $\mathrm{proj}_{\|\cdot\|_{1}\leq 1}$ in (4) are applied column-wise to the matrix input. Hereby we summarize MagR for processing one linear layer in Algorithm 1 with the column-wise $\ell_{1}$-ball projection as detailed in Algorithm 2, which generalizes Algorithm 3 in Appendix A.1, by handling matrix inputs (or batches of vectors).
@@
Input: Pre-trained weight matrix $\hat{\boldsymbol{W}}\in\mathbb{R}^{m\times n}$; Hessian matrix $\boldsymbol{H}=\boldsymbol{X}^{\top}\boldsymbol{X}\in\mathbb{R}^{m\times m}$; max iteration number $K$; step size $\eta=\frac{1}{\lambda_{\max}(\boldsymbol{H})}$; penalty parameter $\alpha>0$.
@@
Extension to Per-Group Quantization. By using more float scaling factors, per-group quantization becomes a preferred strategy for mitigating accuracy loss at extremely low bit-widths. In this approach, a weight vector $\mathbf{w}\in\mathbb{R}^{m}$ is segmented into groups of weights, each containing $d$ elements, with all weights within a group sharing a common scaling factor for quantization. Here, per-group MagR applies an $\ell_{\infty}$ penalty to each vector of grouped weights. Consequently, the $\ell_{1}$-ball projection is independently performed on these vectors, while maintaining the gradient descent step unchanged. We note that the group-wise $\ell_{1}$-ball projection can be easily done using Algorithm 2, with an additional reshaping of the input $\boldsymbol{V}\in\mathbb{R}^{m\times n}$ into $\mathbb{R}^{d\times\left(\frac{m}{d}\cdot n\right)}$.
@@ 5 Experiments
We employed our MagR processing approach on top of the two gradient-free PTQ methods in main text, RTN and OPTQ frantar2022optq , to quantize the LLaMA1 (7B-65B) touvron2023llama and LLaMA2 (7B-70B) touvron2023llama2 model families. In the Appendix A.2, we extend MagR with QuIP chee2024quip (MagR+QuIP) to quantize LLaMA2 (7B-70B) model families. By applying MagR on top of RTN (MagR+RTN), we achieved better results than AWQ lin2023awq for per-channel INT3 and INT4 weight quantization. Additionally, MagR combined with OPTQ (MagR+OPTQ) achieved state-of-the-art performance for INT3 and INT4 qunatization. To enhance the per-channel INT2 quantization, we ran 30 additional iterations of coordinate descent algorithm behdin2023quantease ; zhang2024comq on top of OPTQ, which we denote by MagR+OPTQ. It turns out MagR+OPTQ is superior to both Omniquant shao2023omniquant and QuIP chee2024quip in terms of perplexity (Table 9), and falls just short of QuIP in zero-shot tasks for 13B and 70B models (Table 4). Note that QuIP uses random orthogonal transformations (so-called Incoherence Processing) to process both the weights and features, resulting in 1.5×1.5$\times$1.5 × slower throughput than OPTQ. In contrast, MagR-based method does not introduce any overhead whatsoever compared with OPTQ.
@@
The choice of parameters. To ensure that the MagR-processed layer output $\boldsymbol{X}\boldsymbol{W}$ is faithful to the original $\boldsymbol{X}\hat{\boldsymbol{W}}$, we need to use a tiny penalty parameter $\alpha$ in (2). For per-channel quantization, $\alpha$ was fixed to be $10^{-3}$ in our experiments, but we did find that setting it to a smaller value of $5\times 10^{-4}$ or $10^{-4}$ can sometimes slightly improve the perplexity (with a relative change of $<1\%$ in ppl). Similarly for per-group quantization, we set $\alpha$ to $10^{-4}$, while reducing it to $5\times 10^{-5}$ or $10^{-5}$ could sometimes also slightly improve the perplexity. An ablation study on $\alpha$ is provided in the Appendix A.2.
Furthermore, we used a multiplicative scalar $\beta<1$ to decay the standard quantization step $\delta=\frac{\max(\boldsymbol{w})-\min(\boldsymbol{w})}{2^{b}-1}$ (or equivalently, the quantization scale) of the quantizer. In other words, our $\delta=\beta\cdot\frac{\max(\boldsymbol{w})-\min(\boldsymbol{w})}{2^{b}-1}$. It has been shown in existing works li2016ternary ; rastegari2016xnor that, optimal quantization step for binary or ternary quantization yielding the minimum quantization error is not given by $\frac{\max(\boldsymbol{w})-\min(\boldsymbol{w})}{2^{b}-1}$. Shrinking $\delta$ at low bit-width results in a more clustered quantization grid lattice that fits the weights better, which leads to a smaller overall error. In general, $\beta$ is positively correlated with the bit-width used. For per-channel quantization, the best $\beta\in[0.8,0.85]$ on INT2 quantization, whereas the empirically optimal $\beta$ is around $0.9$ for INT3 quantization. As for INT4, $\beta$ is simply set to 1, that is, we used the standard quantization step. In addition, for per-group quantization, we chose $\beta=0.95$ for both INT2 and INT3 quantization. The ablation study of $\beta$ is in the Appendix A.2. We observed that this refinement on the quantization step $\delta$ significantly improves the performance of the PTQ method. In addition, the iteration number $K$ in Algorithm 1 was set to $150$ across all the experiments.
@@ 5.1 Language Generation
We concentrate our analysis on perplexity-based tasks. The results for the LLaMA2 family with context length of 2048, are elaborated in Table 9, while those for LLaMA1 are provided in Appendix Table 6. As evidenced by the tables, the MagR preprocessing consistently improve the performance of the baselines RTN and OPTQ. Moreover, MagR+OPTQ consistently outperforms most baseline across the LLaMA family models for both per-channel and per-group weight quantization. Particularly, for INT2, MagR+OPTQ performs 30 additional coordinate descent (CD) iterations behdin2023quantease ; zhang2024comq on top of OPTQ to refine the solution, surpassing all baselines.
@@ 5.2 Zero-Shot Tasks
We evaluated the performance of quantized models on several zero-shot tasks. The results are reported in Table 4. Similar to previous observations, the proposed MagR demonstrates superior performance on most models compared to OmniQuant, with a small gap compared to QuIP chee2024quip . Nonetheless, it is reasonable and commendable that our algorithm achieves results close to QuIP without introducing any inference overhead. It is possible to further improve our approach based on the insight behind QuIP chee2024quip — i.e., quantization benefits from incoherent weight and Hessian matrices; see Table 9 for the results in the appendix.
Table 5: The runtime of MagR+RTN, MagR+OPTQ, and MagR+OPTQ on an Nvidia A100 GPU, with comparisons to their vanilla counterparts, namely, RTN and OPTQ.
@@ Appendix A Appendix / supplemental material
A.1 Projection of Vectors Onto $\ell_{1}$-Ball
In this section, we show how to compute the projection onto the unit $\ell_{1}$-Ball. That is, for any fixed $\boldsymbol{v}\in\mathbb{R}^{m}$, we solve the optimization problem:
$\min_{\boldsymbol{x}\in\mathbb{R}^{m}}\;\|\boldsymbol{x}-\boldsymbol{v}\|^{2}\quad\mbox{subject to}\quad\|\boldsymbol{x}\|_{1}\leq 1.$ (5)
Consider the Lagrangian $\mathcal{L}(\boldsymbol{x},\lambda)=\frac{1}{2}\|\boldsymbol{x}-\boldsymbol{v}\|^{2}+\lambda(\|\boldsymbol{x}\|_{1}-1)$, where $\lambda\in\mathbb{R}$ is the lagrange multiplier. Let $\boldsymbol{x}^{*}$ be the optimal solution of (5), then there exits $\lambda^{*}$ such that the following Karush–Kuhn–Tucker (KKT) conditions [6] hold:
  • Stationarity: $\mathbf{0}\in\partial_{\boldsymbol{x}}\mathcal{L}(\boldsymbol{x}^{*},\lambda^{*})\Leftrightarrow\boldsymbol{x}^{*}=\mathrm{sgn}(\boldsymbol{v})\odot\max\{|\boldsymbol{v}|-\lambda^{*},0\}$
  • Slackness: $\lambda^{*}\,(\|\boldsymbol{x}^{*}\|_{1}-1)=0\Leftrightarrow$ $\lambda^{*}=0$ or $\|\boldsymbol{x}^{*}\|_{1}=1$.
  • Primal feasibility: $\|\boldsymbol{x}^{*}\|_{1}-1\leq 0\Leftrightarrow\|\boldsymbol{x}^{*}\|_{1}\leq 1$
  • Dual feasibility: $\lambda^{*}\geq 0\Leftrightarrow$ $\lambda^{*}=0$ or $\lambda^{*}>0$
where $\mathrm{sgn}$ is the signum function applied element-wise on vectors, i.e.,
$\mathrm{sgn}(\boldsymbol{v})_{i}=\begin{cases}1&\mbox{if }v_{i}>0,\\ 0&\mbox{if }v_{i}=0,\\ -1&\mbox{if }v_{i}<0.\end{cases}$
We note that the stationarity condition:
$\boldsymbol{x}^{*}=\mathrm{sgn}(\boldsymbol{v})\odot\max\{|\boldsymbol{v}|-\lambda^{*},0\}$
gives the projection of $\boldsymbol{v}$ onto the unit $\ell_{1}$ ball, provided $\lambda^{*}$ is known. Therefore, what remains is to find $\lambda^{*}$:
  • Case I: $\lambda^{*}=0$. We have $\boldsymbol{x}^{*}=\boldsymbol{v}$, which corresponds to the case that $\|\boldsymbol{v}\|_{1}\leq 1$.
  • Case II: $\lambda^{*}>0$. The slackness condition yields $\|\boldsymbol{x}^{*}\|_{1}=1$, $i.e.$, $\sum_{i=1}^{n}|x_{i}^{*}|=1$, or equivalently, $\sum_{i=1}^{n}\left|\mathrm{sgn}(v_{i})(|v_{i}|-\lambda^{*})_{+}\right|=\sum_{i=1}^{n}(|v_{i}|-\lambda^{*})_{+}=1$ with $x_{+}:=\max\{x,0\}$. That is, $\lambda^{*}$ is the root of the piece-wise linear equation:
    $\sum_{i=1}^{n}(|v_{i}|-\lambda)_{+}=1,$ (6)
    which can be solved by sorting.
In summary, Algorithm 3 details the implementation of projecting $\boldsymbol{v}\in\mathbb{R}^{m}$ onto a general $\ell_{1}$-ball with radius $\epsilon$. In Step 4, we specifically compute the root $\lambda^{*}$ (or $\theta$) of (6) for Case II.
Algorithm 3 Projection onto $\ell_{1}$-ball.
@@ A.1 Projection of Vectors Onto $\ell_{1}$-Ball
1:  if $\|\boldsymbol{v}\|_{1}>\epsilon$ then
2:     Sort $|\boldsymbol{v}|$ into $\boldsymbol{\mu}$ such that $\mu_{1}\geq\mu_{2}\geq\ldots\geq\mu_{m}$.
@@ A.2 Additional Experimental Results
A.3 Ablation Study
Impact of the parameter $\alpha$. The tiny penalty parameter $\alpha$ balances the trade-off between output discrepancy and the maximum magnitude of the weights. We carry out experiments on channel-wise quantization for differernt $\alpha$ on LLaMA2-7B. The choice of $\alpha$ is independent of the bit-width. As shown in Table 7, we can find that both too large and too small $\alpha$ will lead performance degeneration. Compared to INT4, fluctuations in alpha at INT3 result in greater performance fluctuations. Fortunately, $\alpha=0.001$ works well for all channel-wise quantization.
Table 7: The perplexity of quantized LLaMa2-7B models for different $\alpha$ values.
$\alpha$ W/A Wikitext2 (PPL) C4 (PPL)
0.005 4/16 5.84 7.55
0.001 4/16 5.70 7.28
0.0005 4/16 5.72 7.29
0.0001 4/16 5.78 7.35
0.00001 4/16 5.81 7.40
0.005 3/16 6.64 8.74
0.001 3/16 6.41 8.23
0.0005 3/16 6.49 8.38
0.0001 3/16 6.83 8.79
0.00001 3/16 7.08 9.19
Impact of the parameter $\beta$. We shrink the quantization step to reduce the overall quantization error by a multiplicative scalar $\beta$. To investigate the influence of $\beta$, we experiment with different value of $\beta$ at INT3 and INT2 channel-wise quantization. As shown in Table 8, $\beta$ is positively correlated with the bit-width. Specifically, the best $\beta$ is around 0.9 for INT3 quantization and for INT2 quantization the optimal $\beta$ is around 0.8.
Table 8: The perplexity of quantized LLaMa2-7B models for different $\beta$ values.
$\beta$ W/A Wikitext2 (PPL) C4 (PPL)
1 3/16 6.43 8.33
0.9 3/16 6.41 8.23
0.85 3/16 6.48 8.39
0.8 3/16 7.08 9.19
1 2/16 16.99 24.12
0.9 2/16 20.88 31.78
0.85 2/16 16.76 24.45
0.8 2/16 16.73 23.73
Impact of MagR on Quantization Error. To explore how MagR affects quantization error, we compared the errors with and without MagR by randomly select five layers from LLaMA2 models. As illustrated in Figure 2, quantization error is notably reduced across all layers with the application of MagR.
Figure 2: Layer-wise quantization errors (root mse) for MagR+OPTQ and OPTQ, respectively, for 4-bit quantization. The layers are selected randomly for visualization, but improvement is consistent across all layers.
The adaptive capacity of MagR. We investigated the combined effects of MagR and QuIP. As illustrated in Table 9, incorporating MagR significantly enhances the performance of QuIP, leading to improved quantization results for the LLaMA2 models family.
Table 9: Perplexity of MagR+QuIP for LLaMA2 models on Wikitext2 and C4.
Datasets Wikitext2 C4
[…]
FP16 Baseline 5.47 4.88 6.97 6.46
W2A16 QuIP 27.13 10.09 31.33 13.13
MagR+QuIP 13.31 9.40 14.49 11.07
W3A16 QuIP 6.50 5.34 8.74 7.34
MagR+QuIP 6.25 5.29 7.88 7.02
W4A16 QuIP 5.94 5.01 8.01 6.88
MagR+QuIP 5.74 4.99 7.25 6.63